bob is two years older than jim. if the sum of their ages is 104 years. how old is bob?

Answer by stanbon(75887)

bob is two years older than jim. if the sum of their ages is 104 years. how old is bob?
----
Equations:
b = j + 2
b + j = 104
----
Substitute for "b" and solve for "j":
j+2 + j = 104
2j = 102
j = 51 (jim's age now)
Solve for "b":
b = j+2
b = 53 (bob's age now)
==============
Cheers,
Stan H.